# Key varieties for prime $\mathbb{Q}$-Fano threefolds defined by Freudenthal triple systems

### v2

@@ 1.1. Classification of $\mathbb{Q}$-Fano threefolds
A complex projective variety is called a $\mathbb{Q}$-Fano variety if it is a normal variety with only terminal singularities, and its anti-canonical divisor is ample. A $\mathbb{Q}$-Fano variety is called prime if its anti-canonical divisor generates the group of numerical equivalence classes of $\mathbb{Q}$-Cartier divisors. This paper concerns with the classification of prime $\mathbb{Q}$-Fano 3-folds and is a companion paper to [Tak3, Tak4, Tak6, Tak7], where we construct certain affine varieties and show that they produce, as weighted complete intersections of their appropriate weighted projectivizations, several examples of prime $\mathbb{Q}$-Fano 3-folds. We call such affine varieties or its appropriate weighted projectivizations key varieties for $\mathbb{Q}$-Fano $3$-folds. The prime $\mathbb{Q}$-Fano 3-folds constructed in ibid. are of anti-canonical codimension 4, where, by the anti-canonical codimension of a $\mathbb{Q}$-Fano 3-fold $X$, we mean the codimension of $X$ in the weighted projective space of the minimal dimension determined by the anti-canonical graded ring of $X$.
The affine varieties constructed in ibid. are related with $\mathbb{P}^{2}\times\mathbb{P}^{2}$-fibration. In this paper, we construct several affine varieties related with $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibration, and produce, as appropriate weighted complete intersections of their weighted projectivizations, certain examples of prime $\mathbb{Q}$-Fano 3-folds of anti-canonical codimension 4. Ahead of our study including ibid., Coughlan and Ducat did similar work in [CD] defining the $C_{2}$- and $G_{2}^{(4)}$-cluster varieties and using them as key varieties, where the former is related with $\mathbb{P}^{2}\times\mathbb{P}^{2}$-fibration, and the latter is related with $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibration.
Actually, further ahead of [CD] and ibid., Papadakis [P1, P2] constructed via the theory of unprojection more general affine varieties related with $\mathbb{P}^{2}\times\mathbb{P}^{2}$- or $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibration with nicknames Tom and Jerry respectively. They seem, however, experimentally too large as key varieties for prime $\mathbb{Q}$-Fano $3$-folds; in [CD] and ibid., to produce prime $\mathbb{Q}$-Fano 3-fold, we extract appropriate subvarieties from Papadakis’ affine varieties. It is non-trivial what kind of subvarieties are chosen. The cluster varieties was discovered in [CD] via mirror symmetry of log Calabi-Yau surfaces. One purpose of this paper is to reveal that the theory of Freudenthal triple system is a natural framework to define our key varieties and to describe their equations.
@@ 1.2. Freudenthal triple system
In 1960s and 70s, axiomatic definitions of FTS were given in several researches, which start from a vector space with a symplectic form and a tri-linear product and a quartic form defined from them. Among such researches, we basically follow Ferrar’s one [Fe] in this paper; the definition and the way of investigation of FTS given by him turn out to be quite suitable to describe our key varieties conceptually.
@@ 1.3. Key varieties for prime $\mathbb{Q}$-Fano threefolds
For a general 8-dimensional FTS over the complex number field $\mathbb{C}$, the locus consisting of strictly regular elements is isomorphic to the affine cone over $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$ (Corollary 2.31). From our coordinatization of an 8-dimensional FTS mentioned above, we may define an affine scheme $\mathfrak{F}_{\mathbb{A}}^{22}$ related with $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibration (Definition 3.1). It is a subscheme of Papadakis’ Jerry as mentioned above but is still too large to produce prime $\mathbb{Q}$-Fano 3-folds. We extract several subvarieties of $\mathfrak{F}_{\mathbb{A}}^{22}$ and obtain several results as for key varieties of prime $\mathbb{Q}$-Fano 3-folds. Among other things, we state the following as the main result of this paper, which summarizes Propositions 4.4, 4.5, 4.10, and Theorem 4.11:
@@ 1.4. Structure of the paper
In Part 2, we first construct an affine scheme $\mathfrak{F}_{\mathbb{A}}^{22}$ from a natural parameterization of an 8-dimensional FTS. In Sections 46, we extract three subvarieties $\mathfrak{U}_{\mathbb{A}}^{14}$ ,$\mathfrak{S}_{\mathbb{A}}^{8}$, $\mathfrak{Z}_{\mathbb{A}}^{12}$ respectively of $\mathfrak{F}_{\mathbb{A}}^{22}$ and show that they produce, as weighted complete intersections of their weighted projectivizations, certain prime $\mathbb{Q}$-Fano 3-folds. Several properties of $\mathfrak{U}_{\mathbb{A}}^{14}$ ,$\mathfrak{S}_{\mathbb{A}}^{8}$, $\mathfrak{Z}_{\mathbb{A}}^{12}$ are also obtained and they are important to construct prime $\mathbb{Q}$-Fano 3-folds. We describe $\mathfrak{U}_{\mathbb{A}}^{14}$ in detail in Section 4 (we refer to Theorem 1.1 for a summary). As is clarified in this paper, the affine variety $\mathfrak{U}_{\mathbb{A}}^{14}$ is also the cornerstone for studying other key varieties. Moreover, it will produce other prime $\mathbb{Q}$-Fano 3-folds in our future work. By these reasons, we devote many pages to the investigation of $\mathfrak{U}_{\mathbb{A}}^{14}$. As for the variety $\mathfrak{S}_{\mathbb{A}}^{8}$, we describe in detail its equations, an ${\rm SL}_{2}\rtimes{\rm SL}_{2}$-action on it, and its relation with $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibration. We revisit $\mathfrak{S}_{\mathbb{A}}^{8}$ in Section 7. As for the variety $\mathfrak{Z}_{\mathbb{A}}^{12}$, we show in Section 6 that a certain weighted projectivization of it is the $11$-dimensional $\mathbb{Q}$-Fano variety constructed in the paper [Tak5].
In Section 7, we revisit the $G_{2}^{(4)}$-cluster variety defined in [CD], which we denote by $\mathfrak{Cl}_{\mathbb{A}}^{10}$. By the big table [Table] of prime $\mathbb{Q}$-Fano 3-folds obtained from $\mathfrak{Cl}_{\mathbb{A}}^{10}$, we observe that such prime $\mathbb{Q}$-Fano 3-folds are weighted complete intersections of some weighted projectivizations of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ itself or its several subvarieties. In Subsections 7.37.7, we show that $\mathfrak{Cl}_{\mathbb{A}}^{10}$ itself and such subvarieties are actually isomorphic to subvarieties of $\mathfrak{U}_{\mathbb{A}}^{14}$ or $\mathfrak{Z}_{\mathbb{A}}^{12}$ weighted homogeneously with respect to some weights of coordinates. Hence it turns out that all prime $\mathbb{Q}$-Fano 3-folds obtained from $\mathfrak{Cl}_{\mathbb{A}}^{10}$ are also obtained from $\mathfrak{U}_{\mathbb{A}}^{14}$ or $\mathfrak{Z}_{\mathbb{A}}^{12}$. Among other things, we also show that the subvariety of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ studied in Subsection 7.4 is weighted homogeneously isomorphic to $\mathfrak{S}_{\mathbb{A}}^{8}$, and the subvarieties $\mathfrak{T}_{\mathbb{A}}^{8}$ and $\mathfrak{B}_{\mathbb{A}}^{6}$ of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ studied in Subsections 7.5 and 7.7 admit a nontrivial ${\rm SL}_{2}$-action.
@@ 1.5. A future plan
As is noted in Subsection 1.4, the affine variety $\mathfrak{U}_{\mathbb{A}}^{14}$ produces examples of prime $\mathbb{Q}$-Fano $3$-folds which are also obtained from the $G_{2}^{(4)}$-cluster variety $\mathfrak{Cl}_{\mathbb{A}}^{10}$. By Theorem 1.1 (2), one more example is already added in this paper. Actually, we can verify that the affine variety $\mathfrak{U}_{\mathbb{A}}^{14}$ produce more examples of prime $\mathbb{Q}$-Fano 3-folds. Since the verification takes more pages (cf. [Tak3, Tak4]), we will publish it elsewhere.
@@
  • $w(*)$: the weight of coordinate * of a weighted projective space.
@@
    $\begin{pmatrix}\begin{vmatrix}f_{2}&f_{3}\\ g_{2}&g_{3}\end{vmatrix}&-\begin{vmatrix}f_{1}&f_{3}\\ g_{1}&g_{3}\end{vmatrix}&\begin{vmatrix}f_{1}&f_{2}\\ g_{1}&g_{2}\end{vmatrix}\end{pmatrix}.$
@@
Acknowledgment. I am grateful to Professor Shigeru Mukai; inspired by his articles [Mu1, Mu2], I was led to Jordan algebra and FTS to describe key varieties. Moreover, I was helped to find the key variety $\mathfrak{U}_{\mathbb{A}}^{14}$ by his inference as for dimensions of key varieties. I would like to dedicate this paper to him on his 70th birthday. This work is supported in part by Grant-in Aid for Scientific Research (C) 16K05090.
@@ 2. Strictly regular elements in Freudenthal triple systems
Throughout this section, we assume that $\mathsf{k}$ is a field of characteristic $2,3$.
@@ 2.2. Jordan algebraic description of FTS
Assume that $N_{x}(x)$ has a multiple factor $l(x)$, which must be a linear form since $N_{x}(x)$ is a cubic form. We choose coordinates $x_{1},\dots,x_{n}$ of $V_{x}$ such that $l(x)=x_{1}$ and $N_{x}(x)=x_{1}^{2}x_{2}$ (recall that $n\geq 2).$ Then, for $a={}^{t}\!(a_{1},\dots,a_{n}),$ $b={}^{t}\!(b_{1},\dots,b_{n})\in V_{x}$ (we consider $a,b$ as column vectors), we have $\partial_{a}N_{x}(b)=2a_{1}b_{1}b_{2}+a_{2}b_{1}^{2}$, which is equal to $\beta(a,b^{\sharp})$ by Corollary 2.8. Note that the subset $\{b^{\sharp}\mid b\in V_{x}\}\subset V_{y}$ is dense in $V_{y}$ by Proposition 2.10 and the assumption that $N_{x}(x)$ is not identically zero. Therefore, by the assumption that $n\geq 3$, the equality $\beta(a,b^{\sharp})=2a_{1}b_{1}b_{2}+a_{2}b_{1}^{2}$ implies that $\beta$ is degenerate , a contradiction to (2.4).
@@
(2.12) $x\bullet x^{\sharp}\bullet\widetilde{x}=1/6\,(\beta(\widetilde{x},x^{\sharp})x-N_{x}(x)\widetilde{x}),$
@@ 2.3. Condition on strict regularity
$+\left(st-1/3\,\beta(x,y)\right)\widetilde{x}+2\widetilde{t}\left(sx-y^{\sharp}\right)-2\widetilde{y}\sharp(ty-x^{\sharp})+2\Delta_{x}(x,y;\widetilde{x})$
$-\left(st-1/3\,\beta(x,y)\right)\widetilde{y}-2\widetilde{s}\left(ty-x^{\sharp}\right)+2\widetilde{x}\sharp(sx-y^{\sharp})-2\Delta_{y}(x,y;\widetilde{y}),$
@@
(2.19) $\beta\left(\Delta_{x}(x,y;\widetilde{x}),\widetilde{y}\right)=\beta\left(\widetilde{x},\Delta_{y}(x,y;\widetilde{y})\right).$
@@
$\beta\left(\Delta_{x}(x,y;\widetilde{x}),\widetilde{y}\right)$
@@
We denote by $\mathfrak{R}$ the affine scheme in $V$ defined by the equations (2.21), (2.22), and (2.23) as in Proposition 2.18.
@@
Here we mention a few background of Theorem 2.18. Our derivation of the defining equations of $\mathfrak{R}$ as in Theorem 2.18 is inspired by [Kr, Lem.23] and [YamA, p.253–254], and its origin is traced back to the series of the fundamental papers [Fr] (see also [Cl, Prop.6.2] and [Fa, (2.1)]). We refer to [Cl, Prop.6.2] and [KaYas, p.515, Section 0] for aspects of this equation of $\mathfrak{R}$ in a graded Lie algebra of contact type. In [KaYas], the projectivization of $\mathfrak{R}$ is called Freudenthal variety. In [Mu1, Section 5] and [Mu2], the projectivization of $\mathfrak{R}$ is studied in relation with smooth Fano threefolds and is called Legendre projective variety.
@@ 2.4. More on $\Delta_{x}$ and $\Delta_{y}$
In this subsection, we examine the condition (2.23) more in detail. It turns out that we need this condition for several $\widetilde{x}\in V_{x}$ in general to define $\mathfrak{R}$.
@@
By (2.20) as in Lemma 2.16, we have $\beta\left(\Delta_{x}(x,y;\widetilde{x}),\widetilde{y}\right)=\beta\left(x,\Delta_{y}(\widetilde{x},y;\widetilde{y})\right)$. Therefore, (1) is equivalent to that $\beta\left(x,\Delta_{y}(\widetilde{x},y;\widetilde{y})\right)=0$ for any $x\in V_{x}$ and $y\in V_{y}$. By nondegeneracy of $\beta$, this is also equivalent to
@@
We show (1) implies (2). Assume that (1) holds. We set $y=\widetilde{x}^{\sharp}$ in the equality (2.25). Then we have $\beta(\widetilde{x},y)=\beta(\widetilde{x},\widetilde{x}^{\sharp})=3N_{x}(\widetilde{x})$ by (2.9). By (2.13), we also have
$\widetilde{x}\bullet y\bullet\widetilde{y}=\widetilde{x}\bullet\widetilde{x}^{\sharp}\bullet\widetilde{y}=1/6(-\beta(\widetilde{x},\widetilde{y})\widetilde{x}^{\sharp}+N_{x}(\widetilde{x})\widetilde{y}).$
@@
Now we show (2) implies (1). Assume that (2) holds. We write $\widetilde{y}=\alpha\widetilde{x}^{\sharp}$ for some $\alpha\in\mathsf{k}$ (actually, computing $\beta(\widetilde{x},\widetilde{y}),$ we have $\alpha=\frac{\beta(\widetilde{x},\widetilde{y})}{3N_{x}(\widetilde{x})}$). Inserting $\widetilde{y}=\alpha\widetilde{x}^{\sharp}$ in the l.h.s. of (2.25), we see that the equality (2.25) holds for any $y\in V_{y}$ by Lemma 2.21. ∎
@@
generate an $(n-1)$-dimensional vector space.
@@
    Let $V$ be a $3$-dimensional vector space with coordinates $x_{1},x_{2},x_{3}$. The cubic form $x_{1}x_{2}x_{3}$ define a Jordan algebra structure on $V$, whose $\sharp$-mapping is $V\ni(x_{1},x_{2},x_{3})\mapsto(x_{2}x_{3},x_{1}x_{3},x_{1}x_{2})\in V,$ and the bi-linear trace is $\beta(x,y)=x_{1}y_{1}+x_{2}y_{2}+x_{3}y_{3}$ for $x=(x_{1},x_{2},x_{3})$ and $y=(y_{1},y_{2},y_{3}).$ By an explicit calculation, we see that the condition (2.23) is reduced to the two equations:  $0\,2x_{1}y_{1}-x_{2}y_{2}-x_{3}y_{3}=0,x_{1}y_{1}-2x_{2}y_{2}+x_{3}y_{3}=$ 0 . Then we can write down the 9 equations of $\mathfrak{R}$ as follows:
@@
    The space of $3\times 3$ matrix $M(3,\mathsf{k})$ is a Jordan algebra of a cubic form. The Jordan product $\cdot_{J}$ is defined as follows:
@@
    The associated cubic form $N$ is the determinant of a matrix, and for $X\in M(3,\mathsf{k}),$ $X^{\sharp}$ is the adjoint matrix. The bi-linear trace $\beta$ is defined as follows:
@@
    The space of $3\times 3$ symmetric matrix ${\rm Sym}(3,\mathsf{k})$ is also a Jordan algebra of a cubic form, and is a Jordan subalgebra of $M(3,\mathsf{k})$. In this setting, $\mathfrak{R}$ is the $6$-dimensional symplectic Grassmannian ${\rm Sp}(3,6)$ and we can compute the defining equation of $\mathfrak{R}$ as follows:
@@
By Example 2.24 (2) and (3), we observe that, for the defining equations of $\mathfrak{R}$, we need $\Delta_{x}(x,y;\widetilde{x})$ for several $\widetilde{x}$ in general (hence we realize that the equations given in [Cl, Prop.6.2] and [Fa, (2.1)] are not sufficient).

### v3

@@ 1.1. Classification of $\mathbb{Q}$-Fano threefolds
A complex projective variety is called a $\mathbb{Q}$-Fano variety if it is a normal variety with only terminal singularities, and its anti-canonical divisor is ample. A $\mathbb{Q}$-Fano variety is called prime if its anti-canonical divisor generates the group of numerical equivalence classes of $\mathbb{Q}$-Cartier divisors. This paper concerns with the classification of prime $\mathbb{Q}$-Fano 3-folds and is a companion paper to [Tak4, Tak5, Tak6, Tak7], where we construct certain affine varieties and show that they produce, as weighted complete intersections of their appropriate weighted projectivizations, several examples of prime $\mathbb{Q}$-Fano 3-folds. We call such affine varieties or its appropriate weighted projectivizations key varieties for $\mathbb{Q}$-Fano $3$-folds. The prime $\mathbb{Q}$-Fano 3-folds constructed in ibid. are of anti-canonical codimension 4, where, by the anti-canonical codimension of a $\mathbb{Q}$-Fano 3-fold $X$, we mean the codimension of $X$ in the weighted projective space of the minimal dimension determined by the anti-canonical graded ring of $X$.
The affine varieties constructed in ibid. are related with $\mathbb{P}^{2}\times\mathbb{P}^{2}$-fibrations. In this paper, we construct several affine varieties related with $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibrations, and produce, as appropriate weighted complete intersections of their weighted projectivizations, certain examples of prime $\mathbb{Q}$-Fano 3-folds of anti-canonical codimension 4. Ahead of our study including ibid., Coughlan and Ducat did similar work in [CD] defining the $C_{2}$- and $G_{2}^{(4)}$-cluster varieties and using them as key varieties, where the former is related with a $\mathbb{P}^{2}\times\mathbb{P}^{2}$-fibration, and the latter is related with a $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibration.
Actually, further ahead of [CD] and ibid., Papadakis [P1, P2] constructed via the theory of unprojection more general affine varieties related with $\mathbb{P}^{2}\times\mathbb{P}^{2}$- or $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibrations with nicknames Tom and Jerry respectively. They seem, however, experimentally too large as key varieties for prime $\mathbb{Q}$-Fano $3$-folds; in [CD] and ibid., to produce prime $\mathbb{Q}$-Fano 3-fold, we extract appropriate subvarieties from Papadakis’ affine varieties. It is non-trivial what kind of subvarieties are chosen. The cluster varieties were discovered in [CD] via mirror symmetry of log Calabi-Yau surfaces. One purpose of this paper is to reveal that the theory of Freudenthal triple system is a natural framework to define our key varieties and to describe their equations.
@@ 1.2. Freudenthal triple system
In 1960s and 70s, axiomatic definitions of FTS were given by several researches, which start from a vector space with a symplectic form and a tri-linear product and then a quartic form defined from them. Among such researches, we basically follow Ferrar’s one [Fe] in this paper; the definition and the way of investigation of FTS given by him turn out to be quite suitable to describe our key varieties conceptually.
@@ 1.3. Key varieties for prime $\mathbb{Q}$-Fano threefolds
For a general 8-dimensional FTS over the complex number field $\mathbb{C}$, the locus consisting of strictly regular elements is isomorphic to the affine cone over $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$ (Corollary 2.31). From our coordinatization of an 8-dimensional FTS mentioned above, we may define an affine scheme $\mathfrak{F}_{\mathbb{A}}^{22}$ related with a $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibration (Definition 3.1). It is a subscheme of Papadakis’ Jerry as mentioned above but is still too large to produce prime $\mathbb{Q}$-Fano 3-folds. We extract several subvarieties of $\mathfrak{F}_{\mathbb{A}}^{22}$ and obtain several results as for key varieties of prime $\mathbb{Q}$-Fano 3-folds. Among other things, we state the following as the main result of this paper, which summarizes Propositions 4.4, 4.5, 4.10, and Theorem 4.11:
@@ 1.4. Structure of the paper
In Part 2, we first construct an affine scheme $\mathfrak{F}_{\mathbb{A}}^{22}$ from a natural parameterization of an 8-dimensional FTS. In Sections 46, we extract three subvarieties $\mathfrak{U}_{\mathbb{A}}^{14}$ $\mathfrak{S}_{\mathbb{A}}^{8}$, $\mathfrak{Z}_{\mathbb{A}}^{12}$ respectively of $\mathfrak{F}_{\mathbb{A}}^{22}$ and show that they produce, as weighted complete intersections of their weighted projectivizations, certain prime $\mathbb{Q}$-Fano 3-folds. Several properties of $\mathfrak{U}_{\mathbb{A}}^{14}$ $\mathfrak{S}_{\mathbb{A}}^{8}$, $\mathfrak{Z}_{\mathbb{A}}^{12}$ are also obtained and they are important to construct prime $\mathbb{Q}$-Fano 3-folds. We describe $\mathfrak{U}_{\mathbb{A}}^{14}$ in detail in Section 4 (we refer to Theorem 1.1 for a summary). As is clarified in this paper, the affine variety $\mathfrak{U}_{\mathbb{A}}^{14}$ is also the cornerstone for studying other key varieties. Moreover, it will produce other prime $\mathbb{Q}$-Fano 3-folds in our future work. By these reasons, we devote many pages to the investigation of $\mathfrak{U}_{\mathbb{A}}^{14}$. As for the variety $\mathfrak{S}_{\mathbb{A}}^{8}$, we describe in detail its equations, an ${\rm SL}_{2}\rtimes{\rm SL}_{2}$-action on it, and its relation with a $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibration. We revisit $\mathfrak{S}_{\mathbb{A}}^{8}$ in Section 7. As for the variety $\mathfrak{Z}_{\mathbb{A}}^{12}$, we show in Section 6 that a certain weighted projectivization of it is the $11$-dimensional $\mathbb{Q}$-Fano variety constructed in the paper [Tak3].
In Section 7, we revisit the $G_{2}^{(4)}$-cluster variety defined in [CD], which we denote by $\mathfrak{Cl}_{\mathbb{A}}^{10}$. By the big table [Table] of prime $\mathbb{Q}$-Fano 3-folds obtained from $\mathfrak{Cl}_{\mathbb{A}}^{10}$, we observe that such prime $\mathbb{Q}$-Fano 3-folds are weighted complete intersections of some weighted projectivizations of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ itself or its several subvarieties. In Subsections 7.37.7, we show that $\mathfrak{Cl}_{\mathbb{A}}^{10}$ itself and such subvarieties are actually isomorphic to subvarieties of $\mathfrak{U}_{\mathbb{A}}^{14}$ or $\mathfrak{Z}_{\mathbb{A}}^{12}$ weighted homogeneously with respect to some weights on the coordinates. Hence it turns out that all prime $\mathbb{Q}$-Fano 3-folds obtained from $\mathfrak{Cl}_{\mathbb{A}}^{10}$ are also obtained from $\mathfrak{U}_{\mathbb{A}}^{14}$ or $\mathfrak{Z}_{\mathbb{A}}^{12}$. Among other things, we also show that the subvariety of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ studied in Subsection 7.4 is weighted homogeneously isomorphic to $\mathfrak{S}_{\mathbb{A}}^{8}$, and the subvarieties $\mathfrak{T}_{\mathbb{A}}^{8}$ and $\mathfrak{B}_{\mathbb{A}}^{6}$ of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ studied in Subsections 7.5 and 7.7 admit a nontrivial ${\rm SL}_{2}$-action.
@@ 1.5. A future plan
As is noted in Subsection 1.4, the affine variety $\mathfrak{U}_{\mathbb{A}}^{14}$ produces examples of prime $\mathbb{Q}$-Fano $3$-folds which are also obtained from the $G_{2}^{(4)}$-cluster variety $\mathfrak{Cl}_{\mathbb{A}}^{10}$. By Theorem 1.1 (2), one more example is already added in this paper. Actually, we can verify that the affine variety $\mathfrak{U}_{\mathbb{A}}^{14}$ produce more examples of prime $\mathbb{Q}$-Fano 3-folds. Since the verification takes more pages (cf. [Tak7]), we will publish it elsewhere.
@@
  • w()w(*): the weight on the coordinate * of a weighted projective space.
@@
    $\wedge^{2}M=\begin{pmatrix}\begin{vmatrix}f_{2}&f_{3}\\ g_{2}&g_{3}\end{vmatrix}&-\begin{vmatrix}f_{1}&f_{3}\\ g_{1}&g_{3}\end{vmatrix}&\begin{vmatrix}f_{1}&f_{2}\\ g_{1}&g_{2}\end{vmatrix}\end{pmatrix}$.
@@
Acknowledgment. I am grateful to Professor Shigeru Mukai; inspired by his articles [Mu1, Mu2], I was led to Jordan algebras and FTSs to describe key varieties. Moreover, I was helped to find the key variety $\mathfrak{U}_{\mathbb{A}}^{14}$ by his inference for the dimensions of key varieties. I would like to dedicate this paper to him on his 70th birthday. This work is supported in part by Grant-in Aid for Scientific Research (C) 16K05090.
@@ 2. Strictly regular elements in Freudenthal triple systems
Throughout this section, we assume that $\mathsf{k}$ is a field of characteristic $\not=2,3$.
@@ 2.2. Jordan algebraic description of FTS
[…]
@@
[…]
@@ 2.3. Condition on strict regularity
[…]
@@
[…]
@@
[…]
@@
We denote by $\mathfrak{R}$ the affine scheme in $V$ defined by the equations (2.21), (2.22), and (2.23) as in Theorem 2.18.
@@
Here we mention some background of Theorem 2.18. Our derivation of the defining equations of $\mathfrak{R}$ as in Theorem 2.18 is inspired by [Kr, Lem.23] and [YamA, p.253–254], and its origin is traced back to the series of the fundamental papers [Fr] (see also [Cl, Prop.6.2] and [Fa, (2.1)]). We refer to [Cl, Prop.6.2] and [KaYas, p.515, Sec.0] for aspects of this equation of $\mathfrak{R}$ in a graded Lie algebra of contact type. In [KaYas], the projectivization of $\mathfrak{R}$ is called Freudenthal variety. In [Mu1, Sec.5] and [Mu2], the projectivization of $\mathfrak{R}$ is studied in relation with smooth Fano threefolds and is called Legendre projective variety.
@@ 2.4. More on $\Delta_{x}$ and $\Delta_{y}$
In this subsection, we examine the condition (2.23) more in detail.
@@
[…]
@@
[…]
@@
[…]
@@
generate an (n1)(n-1)-dimensional vector space.
@@
    Let $V$ be a $3$-dimensional vector space with coordinates $x_{1},x_{2},x_{3}$. The cubic form $x_{1}x_{2}x_{3}$ define a Jordan algebra structure on $V$, whose $\sharp$-mapping is $V\ni(x_{1},x_{2},x_{3})\mapsto(x_{2}x_{3},x_{1}x_{3},x_{1}x_{2})\in V,$ and the bi-linear trace is $\beta(x,y)=x_{1}y_{1}+x_{2}y_{2}+x_{3}y_{3}$ for $x=(x_{1},x_{2},x_{3})$ and $y=(y_{1},y_{2},y_{3}).$ By an explicit calculation, we see that the condition (2.23) is reduced to the two equations:  $2x_{1}y_{1}-x_{2}y_{2}-x_{3}y_{3}=0,x_{1}y_{1}-2x_{2}y_{2}+x_{3}y_{3}=0.$ Then we can write down the 9 equations of $\mathfrak{R}$ as follows:
@@
    The space of $3\times 3$ matrices $M(3,\mathsf{k})$ is a Jordan algebra of a cubic form. The Jordan product $\cdot_{J}$ is defined as follows:
@@
    The associated cubic form $N$ is the determinant of a matrix, and for $\in M(3,\mathsf{k}), XX^{\sharp}$ is the adjoint matrix. The bi-linear trace $\beta$ is defined as follows:
@@
    The space of $3\times 3$ symmetric matrices ${\rm Sym}(3,\mathsf{k})$ is also a Jordan algebra of a cubic form, and is a Jordan subalgebra of $M(3,\mathsf{k})$. In this setting, $\mathfrak{R}$ is the $6$-dimensional symplectic Grassmannian ${\rm Sp}(3,6)$ and we can compute the defining equation of $\mathfrak{R}$ as follows: